OpenStudy (anonymous):

What is the perimeter of the following polygon? Three sided polygon with height measuring 15 and base measuring 16. The height is a perpendicular bisector of the base. a.50 units b.48 units c.47 units d.46 units

OpenStudy (anonymous):

@ivettef365 can you help

OpenStudy (anonymous):

perimeter = ns n is the number of sides s is the length of any side

OpenStudy (anonymous):

@faithrose101 I don't get it still

OpenStudy (ivettef365):

ok that is a triangle, with perpendicular bisector, h= 15

OpenStudy (anonymous):

okay

OpenStudy (ivettef365):

to find the missing side, use Pythagorean Theorem, so you have 8^2 + 15^2 = c^2

OpenStudy (ivettef365):

64 + 225 = c^2 289 = c^2 sqrt289 = sqrt c^2 17 = c

OpenStudy (ivettef365):

so now you have all 3 sides, 17 + 17 + 16

OpenStudy (anonymous):

i got 50 units

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

okay thanks for the help

OpenStudy (ivettef365):

yw :)